OpenStudy (anonymous):

Identity the graph of the equation 2x^2-4xy+2y^2-7x+7y= 0 then find the angle of rotation

OpenStudy (cwrw238):

this can be written as 2(x - y)(x - y) - 7(x - y) = 0

OpenStudy (cwrw238):

so can u simplify this?

OpenStudy (cwrw238):

(x - y)(2x - 2y - 7) = 0 giving y = x and 2x - 2y - 7 = 0 so y = (1/2) (2x - 7)

OpenStudy (cwrw238):

oh - looks like you're not interested any more..

OpenStudy (cwrw238):

i'm showing you how to go about solving it

OpenStudy (cwrw238):

the result y=x shows that its a straight line going through the origin and the other result gives a straight line y = x - 3.5 which is a straight line parallel to x=y but passing through the y-axis at -3.5 you have 2 parallel straight lines here

OpenStudy (cwrw238):